OpenStudy (anonymous):

It takes a stone 8 seconds to hit the ground when dropped from the top of a building. Give your answers correct to one decimal place. How high is the building? (in meters)

OpenStudy (anonymous):

use the equation of constant acceleration and let g = 9.8 m s^-2 s = ?, a=9.8, t = 8. u= 0 s = ut + 0.5* 9.8 * 8^2 s = 0 + 4.9 * 64 = 313.6 m

OpenStudy (anonymous):

Are you sure this isn't a physics problem? anyways: x_f = x_o + v_o*t + 0.5a*t^2 x_o and v_o both equal zero, and acceleration due to gravity is 9.8. Solve for x_f

OpenStudy (anonymous):

h = ut + 1/2 gt^2 h = 0*8 + 0.5*9.8*8^2 = 0 + 313.6 = 313.6 m